### v1

@@ 4. Results
Our workshop’s structured, case-based deliberations yielded nuanced insights into the multifaceted tensions that arise when using LLMs for legal advice. We identified considerations and concerns across our qualitative data, grouping them into two categories: (1) Dimensions capture contextual factors experts considered when determining appropriate AI responses (Section 4.1); (2) Responses cover desired AI response strategies and guiding principles (Section 4.2).
@@ 4.1. Dimensions
We identified 25 key dimensions that impacted experts’ preferences when it comes to appropriate AI responses. We classified dimensions into four categories: (1) user attributes and behaviors, (2) nature of queries, (3) AI capabilities, and (4) social impacts. Figure 3 outlines these four categories. We now describe each dimension in more detail.
Figure 3. 4-dimension Framework. Experts considered 25 dimensions to determine appropriate AI responses, resulting in a 4-dimensional framework inspired by Kim et al. (Kim et al., 2023a)’s “human-AI-context” factors that affect users’ trust in AI systems. The “Query” dimensions focus primarily on legal considerations, while other dimensions have broader applicability across professional domains.
@@ 4.1.1. User Dimensions
Our participants identified 8 user-related dimensions that AI systems should consider that broadly break down into dimensions related to (1) User attributes and (2) User behavior. Regarding user attributes, or information regarding the person making the query, experts specified four key dimensions of interest:
  • Identity and background, like age, nationality, ethnicity, and vulnerable group status. Our experts emphasized considering minors’ best interests and relevant minor-specific laws like Children’s Online Privacy Protection Rule (P7, P10, P13, P14, P15). Also, nationality (P12), ethnicity (P10), immigration status such as “a DACA recipient” (P12) are also worth considering. Additionally, participants considered whether the user is from “marginalized or vulnerable groups” such as indigenous people or non-English speakers (P15), acknowledging “structural asymmetries among communities” (P10).
  • Geographic location. Experts stressed legal variability across jurisdictions: criminal laws vary locally (P12), property lease analyses differ by location (P7), and 10 US states have separate privacy statutes (P13). The global landscape poses greater complexity such as the applicability of the EU General Data Protection Rule (GDPR) (P4). Interpreting Mexican or Columbian laws requires grasping those countries’ unique histories (P10).
  • Legal sophistication. Our experts noted that the sophistication level of the user should guide the nature of AI legal advice. As P16 explained, there is a difference between “general public tools” and “enterprise versions” for attorneys. Since attorneys bear the ultimate legal liability, professionally-oriented AI tools likely pose fewer risks for misuse. More broadly, P20 suggested having the AI provide advanced and detailed advice to sophisticated users, like a corporate client, already familiar with the technology’s limitations.
@@
User behavior dimensions can sometimes be inferred from inputs but are likely not explicitly stated, such as the reliability of the details provided or the intentions behind the query. This category emerged as experts emphasized lawyers cannot blindly accept user-provided facts. Instead, they must actively probe to construct a complete situational picture before rendering advice. Our findings reveal four key behavioral dimensions for AI systems to assess:
  • Ambiguity. Experts stated that if users’ inputs do not provide enough details about the situation, it is either impossible or risky to provide detailed answers as answers are likely to be flawed (P1, P6, P13). P1 noted, “So many facts are missing. I’m so nervous about the idea of the chat [giving] you legal advice.”
@@
  • Intent. Participants also wanted to clearly understand the underlying intent of the users. P13 stated, users “may also just be doing a really bad job of describing [their] situation…[We need to ask] ‘Are you sure you really mean that?”’ Some participants were wary of AI systems being used to serve the user’s malicious intent, such as “to evade law enforcement,” (P20) or “to defend his crime to avoid illegal consequences of their actions.” (P10)
  • Agency. Experts emphasized users’ degree of agency, or whether users are able to act on the legal guidance given. P17 stated, “There’s still consideration beyond giving the advice that someone might still act on that, which is an important consideration.” Unlike medical advice that requires intermediate steps for treatment, the user may have substantial direct “power to take action,” when furnished with legal recommendations (P20).
@@ 4.1.2. Query Dimensions
At its core, legal advice involves applying relevant law to the specific facts of a person’s situation. Our participants identified 9 key dimensions embedded within users’ legal queries that shape what guidance AI systems can provide. We categorized these expert insights around user cases into three interconnected parts: (1) Relevant facts; (2) Relevant laws; and (3) Nature of desired answers.
@@
  • Relevant Facts. Experts emphasized the importance of key facts needed to furnish suitable legal advice. These included granular details around business practices like data collection methods, advertising revenue streams, and the platform’s terms of conditions (P4). Salient contract terms must be clarified, whether in a lease, employment agreement, conflict waiver, or corporate bylaws (P7, P8, P12). It is also essential to have details on additional stakeholders and counter-parties beyond the user such as competitors (P13), victims, or injured parties (P6, P11). In addition, assessing the stakes involved is significant, ranging from financial liability (P16), to loss of work authorizations or deportation (P12), to imprisonment (P11).
@@
  • Relevant laws. Experts emphasized the complexity of many legal issues raised in user cases. Matters involving diverse areas of law (P14) and jurisdictional variation entail complex legal analysis (P4, P12). The evolving legal landscape necessitates constant research, such as IP addresses that are historically considered personally identifiable information but are not treated as such under most state laws (P12). Participants also stressed the unique nature of criminal matters. The heightened risks in prosecution and incarceration, as well as complex human factors in plea bargaining or sentence hearings, make attorney representation essential (P10, P11). Finally, experts pointed to special considerations for subdomains like tax, privacy, and constitutional law as requiring specialized judgment. The tax code is big, complex, and ambiguous and even experienced attorneys make “judgment calls.” (P13, P19) Privacy laws varies substantively state-by-state (P13) and constitutional matters often involve complex values far broader than codified rules (P20). Given complex judgments embedded in these subjects, participants expressed wariness about AI giving advice.
@@
  • Nature of desired answers. Participants stressed that the quality of the answers depends on what the particular user seeks from the conversation. Users may want straightforward informational outputs like when using traditional search engines (P11–13, P16). In this case, presenting the list of relevant laws for users’ further research could be helpful (P12). In contrast, users may expect tailored legal opinions and strategic advice. According to P7, what the user wants out of the answer may include “compliance or optimizing profits, or tax purposes,” or “step by step instructions” based on predictive assessments of outcomes (“Can I win?”). Finally, users may desire additional insight beyond legal matters (P3, 13, P14). P13 noted the potential need to support users emotionally via empathy, support, and acknowledgement. In one case involving a neighbor’s trespassing, P14 suggested an answer involving home protection measures such as dash cams and dogs, not legal recourse.
@@ 4.1.3. AI Capability Dimensions
Participants raised 5 critical dimensions related to the technical capabilities and constraints of state-of-art LLMs. The transient, AI-specific limitations may shift substantially with ongoing advances of research and development, unlike other categories that rely on users’ needs and contexts. Throughout the discussion, experts disagreed at times: some were more optimistic about future development, while others believed that issues like hallucinations might persist.
@@
  • Accuracy. A key concern raised by multiple participants is the accuracy of AI-generated legal information (P1, P3, P7, P8, P11, P13). P1 stressed the evolving nature of law, noting “We don’t know the law changed from yesterday.” P7, P8, and P13 stressed serious hallucination issues that caused an attorney to be sanctioned for citing ChatGPT-generated cases (Weiser, 2023). Only P11 offered a more positive view: “There is a hallucination issue. [But] you could work with a plugin, or a vector database where you had all this stored. If you could do that reliably, that would be a very good user experience.”
@@
  • Context-awareness. Experts questioned LLMs’ capacity to move beyond static recommendations to context-dependent, adaptive guidance tuned to users’ unique constraints and environments (P8, P10–12, P18, P20). As P11 noted, eligibility criteria like demonstrating terminal illness often rely on specific circumstances. Additionally, procedural legal navigation “is not something you can predict by observing…a large data set” (P12). Others critiqued the staleness of training data, arguing that models cannot “address the local context” (P10, P13) as each situation has “idiosyncratic” details” (P18). However, P20 countered that with enough data, models could likely outline standardized advice and steps applicable to most companies.
@@
  • Confidentiality. Experts extensively discussed confidentiality risks (P4, P7–9, P12, P14, P16). Unlike attorney consultations, AI conversations lack privileged protections against discovery in legal proceedings (P9). Users’ admissions of illegal acts in AI conversations could thus become accessible to adversaries or prosecutors. As P12 cautioned, proper warnings are necessary that AI conversations lack confidentiality protections and could be obtained by others with a warrant. Furthermore, some warned against an AI system’s accidental leak of sensitive information (P4).
@@
  • Accountability. Unlike attorneys, AI systems currently sidestep professional accountability for faulty advice (P8, P16–18). While lawyers’ strict code of conduct and negligence liability apply even to informal suggestions (P17), AI systems evade responsibility either through intermediary immunity laws or non-negotiable disclaimer clauses committing users to bear potential damages (P8, P16). Participants emphasized accountability gaps compared to attorney standards that leave users vulnerable if reliant on AI guidance. Given this gap, P18 argued that uncontrolled AI advice effectively constitutes illegally unauthorized practice of law (UPL).
@@
  • Bias. Experts expressed bias concerns that might reproduce structural stereotypes and discrimination (P5, P10, P13, P17, P20). The aggregated data and conversations could slowly skew the AI system’s performance to favor majority demographics unless measures are taken to actively protect minority views (P5). With English-written data predominantly represented in training data, AI responses may skew towards values and perspectives of those populations (P8). Experts emphasized that without deliberate countermeasures, AI recommendations risk perpetuating existing inequities (P13).
@@ 4.1.4. Impact Dimensions
When evaluating appropriate AI responses, participants also considered 2 dimensions of possible ways that responses could have impacts. These impacts covered emotional, ethical, and cultural factors affecting the individual user seeking guidance and impacts to third parties and society more widely.
  • Impact to user. Experts found that AI systems could potentially weigh the possible downsides including what the user may not have considered that could harm them, such as emotional effects or potential consequences in workplace or relationships (P4, P13, P20). P4 emphasized the need for “guardrails” around emotional prompts like questions including self-harm components. P20 noted that what feels morally neutral in one culture may feel problematic in another, especially for minority demographic groups. P13 cautioned that directly influencing users’ emotional states through algorithmic judgment is highly problematic absent oversight, given risks of uncontrolled bias and manipulation.
  • Impact to others. Experts considered “consequences for other people” who are not direct users as a serious concern with regards to AI legal advice systems (P6, P10, P17). This include risks to indirectly affected third parties encompassing direct biases in advice, indirect impacts of how advice gets interpreted into action, and long-term assimilation of values. P6 emphasized unintended consequences for vulnerable groups, using the example of how advice in harassment cases could further victimize previously affected individuals if not carefully designed. Meanwhile, P17 highlighted broader ethical considerations beyond just technically accurate guidance, noting “there is still consideration beyond giving the advice that somebody might still act on that” in ways that create harm despite good intentions of the system.
4.2. AI Responses: Expert-Preferred Response Strategies and Guiding Principles
Our dimensions in Section 4.1 illustrate the complex considerations involved in AI legal assistance. This section uncovers disagreements among experts through a quantitative and qualitative analysis of our workshop data, as we observed varying perspectives on balancing safety, ethics, and helpfulness.
@@ 4.2.1. Quantitative Results
Participants were asked to identify their preferred AI response strategies by choosing one of our 7 provided strategies or producing their own. The loose bell curve distribution in Figure 4, with strategies ranging from the least interactive, content warning and outright refusal, to the most personally-tailored and detailed recommendations, shows that experts preferred informational responses free of opinion. Notably, the least involved template (Content warning) received no votes, and the most involved template (Recommend actions) received few. Further analysis revealed an intriguing relationship regarding experts’ familiarity with AI systems and their receptivity to more tailored and detailed system responses. Regression testing showed a significant positive correlation ($p<0.05$) between their self-reported general AI usage and openness to more customized and detailed output. Further statistical details of our regression test can be found in Appendix C.
Figure 4. Expert-identified Appropriate AI Responses.

### v2

@@ 4. Results
Our workshop’s structured, case-based deliberations yielded nuanced insights into the multi-layered tensions that arise when using LLMs for legal advice. We identified considerations and concerns across our qualitative data, grouping them into two categories: (1) Dimensions capture contextual factors experts considered when determining appropriate LLM responses (Section 4.1); (2) Responses cover desired LLM response strategies and guiding principles (Section 4.2).
@@ 4.1. Dimensions
We identified 25 key dimensions that impacted experts’ preferences for appropriate LLM responses. We classified dimensions into four categories: (1) User attributes and behaviors, (2) Nature of queries, (3) AI capabilities, and (4) Social impacts. Figure 3 outlines these four categories. We now describe each dimension in more detail.
Figure 3. 4-dimension Framework. Experts considered 25 dimensions to determine appropriate LLM responses, resulting in a 4-dimension framework inspired by Kim et al. (Kim et al., 2023b)’s human-AI-context categorization. While our “Query” dimensions focus primarily on legal considerations, other dimensions have broader applicability across various human-AI interactions.
@@ 4.1.1. User Dimensions
Our participants identified 8 user-related dimensions that AI systems should consider that broadly break down into dimensions related to (1) User attributes and (2) User behavior. User attributes include identity and background, geographic location, legal sophistication, and access to resources. These are characteristics that users may explicitly provide or that can be inferred about them. On the other hand, user behavior refers to aspects such as reliability, intent, agency, and ambiguity, which can be deduced from the user’s inputs and interactions with the AI system but are likely not explicitly stated. Regarding user attributes, experts specified four key dimensions:
  • Identity and background, like age, nationality, ethnicity, and vulnerable group status. Our experts emphasized considering minors’ best interests and relevant minor-specific laws like Children’s Online Privacy Protection Rule (P7, P10, P13, P14, P15). Also, nationality (P12), ethnicity (P10), immigration status such as “a DACA recipient” (P12) were perceived to be worth considering. Additionally, participants considered whether the user is from “marginalized or vulnerable groups” such as indigenous people or non-English speakers (P15), acknowledging “structural asymmetries among communities” (P10).
  • Geographic location. Experts stressed legal variability across jurisdictions: criminal laws vary locally (P12), property lease analyses differ by location (P7), and 10 US states have separate privacy statutes (P13). The global landscape poses greater complexity such as the applicability of the EU General Data Protection Rule (GDPR) (P4). Moreover, when interpreting laws from Mexico or Colombia, it is important to consider the unique histories and legal contexts of these countries, which differ from those of the US (P10).
  • Legal sophistication. Our experts noted that the sophistication level of the user should guide the nature of LLM legal advice. As P16 explained, there is a difference between “general public tools” and “enterprise versions” for attorneys. Since attorneys bear the ultimate legal liability, professionally-oriented AI tools likely pose fewer risks for misuse. More broadly, P20 suggested that LLM systems could provide more advanced and detailed advice to sophisticated users, like a corporate client, who are already familiar with the technology’s limitations and are less likely to misinterpret or misuse the information provided.
@@
The user behavior category emerged as experts emphasized that lawyers should not blindly accept user-provided facts. Instead, lawyers must actively probe and ask questions to construct understanding of situation before offering advice. Our findings reveal four key behavioral dimensions for LLM systems to assess:
  • Ambiguity. Experts stated that if user inputs do not provide enough details about the situation, it is either impossible or risky to provide detailed guidance as the LLM outputs are likely to be flawed due to the incomplete information (P1, P6, P13). P1 noted, “So many facts are missing. I’m so nervous about the idea of the chat [giving] you legal advice [based on this incomplete fact].”
@@
  • Intent. Participants also wanted to clearly understand the underlying intent of the users. P13 stated that users may also do a poor job of describing their situation, and the LLM system should ask for clarification by posing questions like “Are you sure you really mean that?” Some participants were wary of LLMs being used to serve the user’s malicious intent, such as “to evade law enforcement,” (P20) or “to defend his crime to avoid illegal consequences of their actions.” (P10)
  • Agency. Experts emphasized users’ degree of agency, or whether users are able to act on the legal guidance given. P17 stated, “There’s still consideration beyond giving the advice that someone might still act on that.” In the legal setting, unlike in medical contexts where treatment requires intermediate steps by professionals, users may have substantial direct “power to take action” when provided with legal recommendations, such as firing an employee or filing a complaint (P20).
@@ 4.1.2. Query Dimensions
Essentially, legal advice involves applying relevant law to the specific facts of a person’s situation. Our participants identified 9 key dimensions embedded within users’ legal queries that shape what guidance AI systems can provide. We categorized these dimensions into three interconnected parts: (1) Relevant facts; (2) Relevant laws; and (3) Nature of desired answers.
@@
  • Relevant Facts. Experts emphasized the importance of key facts needed to furnish suitable legal advice. These included granular details around business practices like data collection methods, advertising revenue streams, and the platform’s terms of conditions (P4). Existing contract terms must be clarified, whether in a lease, employment agreement, conflict waiver, or corporate bylaws (P7, P8, P12). It is also essential to have details on stakeholders and counter-parties such as competitors (P13), victims, or injured parties (P6, P11). In addition, assessing the stakes involved is significant, ranging from financial liability (P16), to loss of work authorizations or deportation (P12), to imprisonment (P11).
@@
  • Relevant laws. Experts underscored the complexity of many legal issues. Matters involving diverse areas of law (P14) and jurisdictional variation involve a complex legal analysis (P4, P12). The evolving legal landscape necessitates constant research. For example, IP addresses were historically considered personally identifiable information but are not treated as such under most state laws (P12). Participants also stressed the unique nature of criminal matters. The heightened risks in prosecution and incarceration, as well as complex human factors in plea bargaining or sentence hearings, make attorney representation essential (P10, P11). P11 exemplified judges’ idiosyncrasies, quoting a religious federal judge in Washington state: “It really helps. If you’re a Christian, and your criminal defendant appearing before him, should always start with a little prayer when you’re doing your sentence hearing.” Experts pointed to special considerations for subdomains like tax, privacy, and constitutional law as requiring specialized judgment. The tax code is big, complex, and ambiguous, so even experienced attorneys should make “judgment calls.” (P13, P19) Privacy laws varies substantively state-by-state (P13) and constitutional matters often involve complex values far broader than codified rules (P20).
@@
  • Nature of desired answers. Participants stressed that the quality of the answers depends on what the particular user seeks from the conversation. Users may want straightforward informational outputs, like when using traditional search engines (P11–13, P16). In this case, presenting the list of relevant laws for users’ further research could be helpful (P12). In contrast, users may expect tailored legal opinions and strategic advice. According to P7, what the user wants out of the answer may include “compliance or optimizing profits, or tax purposes,” or “step-by-step instructions” based on predictive assessments of outcomes (“Can I win?”). Finally, users may desire additional insights beyond legal matters (P3, 13, P14). P13 noted the need to emotionally support users by extending empathy, support, and acknowledgment. In one case involving a neighbor’s trespassing, P14 suggested home protection measures such as installing dashcams and getting dogs, not just legal recourse.
@@ 4.1.3. AI Capability Dimensions
Participants raised 5 critical dimensions related to the technical capabilities and constraints of state-of-art LLMs. The transient, LLM-specific limitations may shift substantially with ongoing advances of research and development, unlike other categories that rely on users’ needs and contexts. Throughout the discussion, experts disagreed at times: some were more optimistic about future development, while others believed that issues like hallucinations might persist.
@@
  • Accuracy. A key concern raised by multiple participants is the accuracy of AI-generated legal information (P1, P3, P7, P8, P11, P13). P1 stressed the evolving nature of law, noting “We don’t know the law changed from yesterday.” P7, P8, and P13 stressed serious hallucination issues that caused a New York attorney to be sanctioned for citing ChatGPT-generated cases (Weiser, 2023). Only P11 offered a more positive view: “There is a hallucination issue. [But] you could work with a plug-in, or a vector database where you had all this stored. If you could do that reliably, that would be a very good user experience.”
@@
  • Context-awareness. Experts questioned LLMs’ capacity to move beyond static recommendations to context-dependent, adaptive guidance tuned to users’ unique constraints and environments (P8, P10–12, P18, P20). As P11 noted, eligibility criteria like demonstrating terminal illness often rely on specific circumstances. Additionally, procedural legal navigation “is not something you can predict by observing…a large data set” (P12). Others critiqued the staleness of training data, arguing that models cannot “address the local context” (P10, P13) as each situation has “idiosyncratic” details (P18). However, P20 countered that with enough data, models could likely outline standardized advice and steps applicable to various types of users.
@@
  • Confidentiality. Experts extensively discussed confidentiality risks (P4, P7–9, P12, P14, P16), which can be differentiated in a practical and normative sense. From a practical perspective, experts warned against an LLM system’s accidental leak of sensitive information (P4), highlighting the potential for unintended breaches of confidentiality. From a normative perspective, unlike attorney consultations, conversations with LLM systems typically lack privileged protections against discovery in legal proceedings (P9). Attorney-client privilege does not extend to communications with third parties, and LLM system providers (e.g., OpenAI) are obligated to produce relevant documents when served with a valid subpoena. Even if an LLM system operates locally, chat records would likely remain unprotected unless a specific rule shields the information from disclosure. As a result, users’ admissions of illegal acts in LLM conversations could thus become accessible to adversaries or prosecutors. P12 cautioned that proper warnings are necessary to inform LLM conversations lack confidentiality protections and could be obtained by others with a court order.
@@
  • Accountability. Unlike attorneys, LLM systems currently sidestep professional accountability for faulty advice (P8, P16–18). While lawyers’ strict code of conduct and negligence liability apply even to informal suggestions (P17), LLM systems evade responsibility either through intermediary immunity laws or non-negotiable disclaimer clauses committing users to bear potential damages (P8, P16). Participants emphasized accountability gaps compared to attorney standards that leave users vulnerable if reliant on LLM guidance. Given this gap, P18 argued that uncontrolled LLM advice effectively constitutes illegally unauthorized practice of law (UPL).
@@
  • Bias. Experts expressed concerns that LLM systems might reproduce structural stereotypes and discrimination (P5, P10, P13, P17, P20). They cautioned that the aggregated data used to train these systems could gradually skew the LLM’s performance to favor majority demographics unless measures are taken to actively protect minority views (P5). Given that English-written data predominantly represented in training datasets, experts noted that LLM responses may disproportionately reflect the values and perspectives of English-speaking populations (P8).
@@ 4.1.4. Impact Dimensions
Experts considered 2 dimensions of possible ways that LLM-generated responses could have on users and society. The first dimension focuses on the individual user seeking guidance, taking into account the emotional, ethical, and cultural factors that may be affected by LLM responses. The second dimension extends beyond the individual, considering the broader impacts on third parties and society as a whole.
  • Impact on users. Experts found that LLM systems could potentially weigh the possible downsides including what the user may not have considered that could harm them, such as emotional effects or potential consequences in workplace or relationships (P4, P13, P20). P4 emphasized the need for “guardrails” around emotional prompts like questions including self-harm components. P13 cautioned that influencing users’ emotional states is highly problematic absent oversight, given risks of uncontrolled bias and manipulation. P20 noted that what feels morally neutral in one culture may feel problematic in another, especially for minority groups.
  • Impact on others. Experts considered “consequences for other people” who are not direct users as a serious concern (P6, P10, P17). These consequences include risks to indirectly affected third parties, such as explicit bias and stereotypes in advice, ensuing impacts of how advice is interpreted and acted upon, and the long-term assimilation of values. P6 emphasized the potential for unintended consequences on vulnerable groups, using the example of how advice in harassment cases could further victimize previously affected individuals. Meanwhile, P17 highlighted broader ethical considerations beyond just technically accurate guidance, including assessing scenarios that create harm despite the good intentions of the advice.
4.2. LLM Responses: Expert-Preferred Response Strategies and Guiding Principles
Our dimensions in Section 4.1 illustrate the complex considerations involved in LLM legal advice. This section uncovers disagreements among experts through a quantitative and qualitative analysis of our workshop data, as we observed varying perspectives on balancing safety, ethics, and helpfulness.
Figure 4. Expert-preferred Response Strategies.
@@ 4.2.1. Quantitative Results
Participants were asked to identify their preferred LLM response strategies by choosing one of our 7 provided strategies or producing their own. The resulting distribution, as shown in Figure 4, resembles a loose bell curve, with strategies ranging from the least interactive (content warning and outright refusal) to the most personally-tailored recommendations. This distribution reveals that experts generally preferred information-focused responses that avoid giving definitive judgment. The strategies at the extremes of spectrum, namely ‘content warning,’ which received no votes, and ‘recommend actions,’ which received few votes, were less favored by the experts. The concentration of votes in the middle of the distribution suggests that experts prioritize providing users with relevant information while refraining from offering explicit recommendations or opinions, striking a balance between assisting users and maintaining the LLM system’s role as an informative tool rather than a decision-maker.
Further analysis revealed an intriguing relationship regarding experts’ familiarity with AI systems and their receptivity to more tailored and detailed system responses. Regression testing showed a significant positive correlation ($p<0.05$) between their self-reported general AI usage and openness to more customized and detailed output. Further statistical details of our regression test can be found in Appendix C.